James Allen Morrow, Curriculum Vitae

Born  September 14, 1941, Little Rock, Arkansas
Education Adamson High School Dallas, Texas, 1959

• B.S., California Institute of Technology, 1963

• Ph.D., Stanford University, 1967.

Awards

• Superior Oil Company Scholarship (1959-63)

• National Science Foundation Fellowship (1963-67)

Positions

• Teaching Assistant, Stanford, (1963-67)

• Instructor in Mathematics, University of California, Berkeley (1967-68)

• Lecturer in Mathematics, University of California, Berkeley (1968-89)

• Assistant Professor of Mathematics, University of Washington, (1969-73)

• Visiting Associate Professor of Mathematics, Rice University, (1973)

• Associate Professor of Mathematics, University of Washington, (1973-78)

• Professor of Mathematics, University of Washington, (1978-present)

• Graduate Program Director, University of Washington, (1983-86)

• Undergraduate Program Director, (1989-93)

• Undergraduate Program Director, (1998- )